40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 14 19:33:04 2009, Alberto, Update, General, Arm Cavity Poles measured again after cleaning the optics last week XarmTF01_fit.pdfYarmTF01_fit.pdfArmCavityFinesseMeasurment.tar.gz
    Reply  Tue Dec 15 05:29:31 2009, Alberto, Update, General, Arm Cavity Poles measured again after cleaning the optics last week 
Message ID: 2418     Entry time: Tue Dec 15 05:29:31 2009     In reply to: 2415
Author: Alberto 
Type: Update 
Category: General 
Subject: Arm Cavity Poles measured again after cleaning the optics last week 

 

 The Y arm cavity pole moved down by 130 Hz, whereas the X arm moved by only 34 Hz. I wonder if that is because Kiwamu spent much more time on cleaning ITMY than on any other optic.

ELOG V3.1.3-